A train leaves a station at 9:00 A.M. traveling at a constant rate of speed. Two hours later a second train, traveling at an average rate of 25 miles per hour more than the first train, leaves the same station, going in the same direction, on a parallel track. If the faster train overtakes the first train at 3:00 P.M. on the same day, what is the average rate of speed of the first train.
Answer by stanbon(29603)

A train leaves a station at 9:00 A.M. traveling at a constant rate of speed. Two hours later a second train, traveling at an average rate of 25 miles per hour more than the first train, leaves the same station, going in the same direction, on a parallel track. If the faster train overtakes the first train at 3:00 P.M. on the same day, what is the average rate of speed of the first train.
-----------
If train two overtakes train one at 3:00 train two must have traveled
from 11:00 to 3:00 or 4 hours. Train one left two hours earlier so
it must have traveled for 6 hours.
---------------
Train one DATA:
Time = 6 hours ; Rate = x mph; Distance = r*t = 6x miles
-----------
Train two DATA:
Time = 4 hurs ; Rate = x+25 mph ; Distance = r*t = 4(x+25)
------------
EQUATION:
distance = distance
6x=4(x+25)
6x=4x+100
2x=100
x=50 mph (speed of 1st train)
--------------
Cheers,
Stan H.